OpenStudy (anonymous):

a chunk of metal weighs 15.45 g is added to 21.27 ml of water. The water rises to 24.78ml calculate density of the metal and explain

OpenStudy (anonymous):

Density=Mass/Volume Mass of metal=15.45 g Volume of metal=(23.78-21.27)ml=2.51ml=2.51cm^3 d=15.45g/2.51cm3 d=6.16g/cm^3

OpenStudy (anonymous):

This is the Archimedes principle in action. The object must displace the water by an amount equal to its own volume (assuming it is completely submerged).